a. You've recently read about a chemical laser that generates a 20.0-cm-diameter, 27.0 MW laser beam. One day, after physics class, you start to wonder if you could use the radiation pressure from this laser beam to launch small payloads into orbit. To see if this might be feasible, you do a quick calculation of the acceleration of a 20.0-cm-diameter, 107 kg, perfectly absorbing block.What speed would such a block have if pushed horizontally 103 m along a frictionless track by such a laser?

b. Unpolarized light of intensity I0 is incident on three polarizing filters. The axis of the first is vertical, that of the second is 40.0 ∘ from vertical, and that of the third is horizontal.What light intensity emerges from the third filter?
Answer #1

a)

Radiation pressure is given by

PR=I/c =P/Ac =(27*106)/(pi*0.22/4)(3*108)=2.8648 N/m2

Force

F=PRA =2.8648*(pi*0.22/4)=0.09 N

acceleration

a=F/m =0.09/107=8.4*10-4m/s2

Speed of the block is

V=sqrt[2ad]=sqrt[2*8.4*10-4*103]

V=0.416 m/s

b)

The light intensity emerges from the third filter is

I3=(Io/2)Cos240Cos2(90-40)

I3=0.12123Io